OpenStudy (anonymous):

f(x) = (2x + 3)(x − 2) and g(x) = 2x2 − 16. Find g(x) − f(x).

OpenStudy (anonymous):

Turnaround, every now and then I get a little bit lonely And you're never coming round Turnaround, every now and then I get a little bit tired Of listening to the sound of my tears Turnaround, every now and then I get a little bit nervous That the best of all the years have gone by Turnaround, every now and then I get a little bit terrified And then I see the look in your eyes Turnaround bright eyes, but every now and then I fall apart Turnaround bright eyes, every now and then I fall apart And I need you now tonight And I need you more than ever And if you only hold me tight We'll be holding on forever And we'll only be making it right 'Cause we'll never be wrong together We can take it to the end of the line Your love is like a shadow on me all of the time (all of the time) I don't know what to do and I'm always in the dark We're living in a powder keg and giving off sparks I really need you tonight Forever's gonna start tonight Forever's gonna start tonight Once upon a time I was falling in love But now I'm only falling apart And there's nothing I can do A total eclipse of the heart Once upon a time there was light in my life But now there's only love in the dark Nothing I can say A total eclipse of the heart Turnaround bright eyes, every now and then I fall apart Turnaround bright eyes, every now and then I fall apart And I need you now tonight (And I need you now) And I need you more than ever And if you only hold me tight (And if you only) We'll be holding on forever And we'll only be making it right ('Cause we'll never be wrong) 'Cause we'll never be wrong together We can take it to the end of the line Your love is like a shadow on me all of the time (all of the time) I don't know what to do and I'm always in the dark We're living in a powder keg and giving off sparks I really need you tonight Forever's gonna start tonight Forever's gonna start tonight Once upon a time I was falling in love But now I'm only falling apart And there's nothing I can say A total eclipse of the heart A total eclipse of the heart A total eclipse of the heart Turn around bright eyes Read more: Bonnie Tyler - Total Eclipse Of The Heart Lyrics | MetroLyrics

OpenStudy (p0sitr0n):

To do it, you need to expand the brackets, and then do pointwise subtractions on same powers of x's. Thus, \[( 2x^2 − 16)-(2x + 3)(x − 2)=2x^2-16-2x^2+4x-3x+6= -10+x\]